An infinite non-conducting sheet has a surface charge density  = 0.10 C/m2 on one side. How far apart are equipotential surfaces whose potentials differ by 50 V?
Answer #1

field due to charged sheet

E = sigma / ( 2 eo)

E = 0.1 * 10^-6 / ( 2* 8.85* 10^-12)

E = 5649.72 N/C

as we know

E = V/d

d = V/E

d = 50 / 5649.72

d = 8.85* 10^-3 m

or

d = 8.85 mm

=====

Comment before rate in case any doubt, will reply for sure.. goodluck